MOA-II Obsevation Log Monthly Summary 2009/10
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 22817-23002 | 187 | 1175 | 70230 |
LMC | 4353-4502 | 151 | 844 | 224520 |
SMC | 869-882 | 15 | 29 | 8060 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 353 | 2048 | 302810 (84h00m410s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
51 | 52 | 133 | 134 | 139 | 9 | 31 | 52 | 112 | 107 | 19 | 25 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
41 | 108 | 22 | 22 | 37 | 35 | 13 | 11 | 13 | 9 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
35 | 34 | 34 | 33 | 33 | 36 | 149 | 147 | 144 | 45 | 37 | 36 |
Field13 | Field14 | Field15 | Field16 |
39 | 42 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
15 | 14 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2009/10/1: I charged helium gas. Pumping. / night time = 34020(sec)
2009/10/2: No obs. / night time = 33840(sec)
2009/10/3: / night time = 37200(sec)
2009/10/4: / night time = 33360(sec)
2009/10/5: I took the LMC fields once with V-band. / night time = 33120(sec)
2009/10/6: / night time = 32940(sec)
2009/10/7: / night time = 32700(sec)
2009/10/8: No obs. / night time = 32520(sec)
2009/10/9: / night time = 32280(sec)
2009/10/10: / night time = 32040(sec)
2009/10/11: I took LMC fields once with V-filter. / night time = 31860(sec)
2009/10/12: / night time = 31560(sec)
2009/10/13: / night time = 31380(sec)
2009/10/14: / night time = 31140(sec)
2009/10/15: / night time = 30900(sec)
2009/10/16: I took LMC fields with V-band. Flat images of V-band (F2992) and R-band (F2995) were taken. / night time = 30660(sec)
2009/10/17: No observation. / night time = 30480(sec)
2009/10/18: / night time = 30240(sec)
2009/10/19: We replaced new helium hoses with the leaking ones. No obs. / night time = 30000(sec)
2009/10/20: Mostly cloudy. GB only. / night time = 29760(sec)
2009/10/21: / night time = 29520(sec)
2009/10/22: We took FLAT images (F3000). / night time = 29280(sec)
2009/10/23: Fine, good seeing / night time = 29040(sec)
2009/10/24: LMC V observation / night time = 28800(sec)
2009/10/25: Vacuum gauge broken / night time = 28560(sec)
2009/10/26: No Observation / night time = 28380(sec)
2009/10/27: Fine, good seeing. / night time = 28140(sec)
2009/10/28: Chip 2 dead / night time = 27900(sec)
2009/10/29: LMC and SMC / night time = 27600(sec)
2009/10/30: Cryostat noise. CCD Temperature high. / night time = 27360(sec)
2009/10/31: Bad weather. No observation. / night time = 27120(sec)
TOTAL NIGHT TIME: 953700(sec)
OPERATING RATE: 40.3 %
GB: 12.3 %
LMC: 27.1 %
SMC: 1.0 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month